# Verb Mirage: Unveiling and Assessing Verb Concept Hallucinations in Multimodal Large Language Models

### v1

@@ Appendix C Details about Experiments on Training-free Hallucination Mitigation
C.1 Experiment Setting
Previous research [leng2024mitigating, chen2424halc] has demonstrated that sampling strategy can affect MLLM performance. Therefore, to evaluate the effect of hallucination mitigation methods fairly, we choose to control the use of decoding strategy. As OPERA [huang2024opera] heavily relies on beam search, we select beam search as the default decoding strategy for evaluation on OPERA and VCD [leng2024mitigating] in the main paper.
As HALC [chen2424halc] is very time-consuming and selects DoLA [chuang2024dola] beam search as the default sampling strategy in their codebase, we select DoLA beam search with beam size 1 as the default sampling strategy in our experiment.
C.2 Additional Experiments on VCD: Sampling as Decoding Strategy
We test VCD with sampling as the decoding strategy as the author claimed in the paper [leng2024mitigating], and get results in Tab.  9. From the result, we can see that VCD also shows inconsistent results. Sometimes sampling+VCD may achieve substantial improvement oversampling only, but the result is not better than beam search only.
@@ C.2 Additional Experiments on VCD: Sampling as Decoding Strategy
Appendix D Failure Case Visualization
@@
Figure 14: Some failure cases of MiniCPM-Llama3-V2.5 [yao2024minicpm].
Figure 15: Some failure cases of Gemini-1.5-Flash [team2023gemini].
Figure 16: Some failure cases of GPT-4-Turbo [achiam2023gpt].

### v3

@@ Appendix C Details about Experiments on Training-free Hallucination Mitigation
Experiment Setting
Previous research (leng2024mitigating) has demonstrated that sampling strategy can affect MLLM performance. Therefore, to evaluate the effect of hallucination mitigation methods fairly, we choose to control the use of decoding strategy. As OPERA (huang2024opera) heavily relies on beam search, we select beam search as the default decoding strategy for evaluation on OPERA, VCD, and other methods in the main paper. All experiments are conducted using a server with 2 A100 40G GPUs.
We test VCD with sampling as the decoding strategy as the author claimed in the paper (leng2024mitigating), and get results in Tab. 8. From the result, we can see that VCD also shows inconsistent results. Sometimes sampling+VCD may achieve substantial improvement over sampling only, but the result is not better than beam search only.
@@
Appendix D Details about Mitigation via Pangea Fine-tuning
Pangea (li2024isolated) gathers many heterogeneous datasets in a unified way. We omit Kinetics-700 (carreira2019short) which makes too large a proportion in Pangea, and HAKE (li2022hake) which is in the same domain as HICO. Among other source datasets, we sample 60K images according to the proportion they make up in Pangea. In this way, we sample massive out-of-domain images, each of which only contains rough action labels.
[…]
We gather MC questions and YN questions together. There are 60K MC questions and 60K YN questions. The ratio of YN questions with correct answers Yes and No is approximately 1:19. The learning rate is 2e-5. Training is conducted using a server with 2 A100 40G GPUs with a total batch size of 64.
Appendix E Failure Case Visualization
@@
Appendix F More Discussion
In this paper, we unveil, assess, and study the verb hallucination problem of MLLMs. Many fields such as human-robot interaction, sports analysis, etc. require deep understanding of action verbs, and have images captured under similar fields of view as input. Therefore, our study can inspire broader scopes.
As the first step, we test many MLLMs as well as outstanding methods targeted object hallucination. In addition, we also tried some ad-hoc methods, such as adding “Pay attention to the verb in question.” to the prompt: the F1 score of LLaVA 1.5 on YN questions is 57.5 while that on YN questions without object references is 62.4, showing hardly any gain. To our best knowledge, verb hallucination is still a much harder problem than object hallucination. As for Qwen2.5-VL-7B, the F1 score of YN and YN without object references are 73.79 and 71.99 respectively. Compared with its F1 score of 85.9 on POPE, its verb hallucination is still severer. We have also tried to analyze the influence from network architectures, but were hindered by the fact that mainstream MLLMs are trained with private data, which renders impossible an apple-to-apple comparison. The field of image generation has seen the emergence of Diffusion models based on noise prediction (rombach2022high) and rectified flow (liu2023flow), and Visual Autoregressive Models (tian2024visual), and we hope similar exploration and wide adoption of paradigm shift in mainstream MLLMs.
Figure 16: Some failure cases of MiniCPM-Llama3-V2.5 (yao2024minicpm).
Figure 17: Some failure cases of Gemini-1.5-Flash (team2023gemini).
Figure 18: Some failure cases of GPT-4-Turbo (achiam2023gpt).